# Zero-Shot Trajectory Planning for Signal Temporal Logic Tasks

### v1

@@ Abstract
Signal Temporal Logic (STL) is a powerful specification language for describing complex temporal behaviors of continuous signals, making it well-suited for high-level robotic task descriptions. However, generating executable plans for STL tasks is challenging, as it requires consideration of the coupling between the task specification and the system dynamics. Existing approaches either follow a model-based setting that explicitly requires knowledge of the system dynamics or adopt a task-oriented data-driven approach to learn plans for specific tasks. In this work, we investigate the problem of generating executable STL plans for systems whose dynamics are unknown a priori. We propose a new planning framework that uses only task-agnostic data during the offline training stage, enabling zero-shot generalization to new STL tasks. Our framework is hierarchical, involving: (i) decomposing the STL task into a set of progress and time constraints, (ii) searching for time-aware waypoints guided by task-agnostic data, and (iii) generating trajectories using a pre-trained safe diffusion model. Simulation results demonstrate the effectiveness of our method indeed in achieving zero-shot generalization to various STL tasks.
@@ 1 Introduction
Signal Temporal Logic (STL) is a formal specification language used to describe the temporal behavior of continuous signals. It has become widely adopted for specifying high-level robotic behaviors due to its expressiveness and the availability of both Boolean and quantitative evaluation measures. Controlling robots under STL task constraints, however, is a challenging problem, as it requires balancing both the satisfaction of the task and the feasibility of the system dynamics. In cases where the environment and system dynamics are fully known, several representative methods have been developed, including optimization-based approaches Raman et al. (2014); Kurtz and Lin (2022); Sun et al. (2022), gradient-based techniques Gilpin et al. (2020); Dawson and Fan (2022), and sampling-based methods Ilyes et al. (2023). However, these methods are often difficult to apply in practical scenarios, where the system dynamics and environment are either unknown or difficult to model.
@@
To address the challenge of unknown dynamics, several learning-based approaches have been proposed. One typical method is reinforcement learning (RL) Aksaray et al. (2016); Balakrishnan and Deshmukh (2019); Kalagarla et al. (2021); Venkataraman et al. (2020); Ikemoto and Ushio (2022); Wang et al. (2024), where an appropriate reward function is designed to approximate the satisfaction of the STL task. However, these methods often struggle with long-horizon STL tasks and lack generalization capabilities across different tasks. Another approach involves first learning a system model and then integrating it with model-based planning methods. For example, in Kapoor et al. (2020), the authors trained a neural network to approximate the system dynamics and combined it with an optimization-based approach. However, this method is limited to simple short-horizon STL tasks due to its high computational cost. In He et al. (2024), the authors used goal-conditioned RL to train multiple goal-conditioned policies, referred to as “skills,” to accomplish specific objectives. They then applied a search algorithm to determine the optimal sequence of “skills” needed to satisfy the given STL tasks. While this approach enables a certain degree of task generalizations, these tasks must be based on pre-defined objectives associated with the skills.
@@
More recently, generative models, such as diffusion models Ho et al. (2020), have emerged as a new approach for generating trajectories for systems with unknown dynamics Janner et al. (2022); Ajay et al. (2022); Chi et al. (2023); Carvalho et al. (2023); Huang et al. (2025), gaining popularity across various applications. Compared to traditional model-based reinforcement learning methods, these generative approaches are better suited for long-horizon decision-making and offer greater test-time flexibility Janner et al. (2022), making them particularly effective for complex tasks. For example, for finite Linear Temporal Logic ($\text{LTL}_{f}$) tasks, Feng et al. (2024a) introduced a classifier-based guidance approach to steer the sampling of diffusion models, ensuring that generated trajectories satisfy $\text{LTL}_{f}$ requirements. Similarly, Feng et al. (2024b) proposed a hierarchical framework that decomposes co-safe LTL tasks into sub-tasks using hierarchical reinforcement learning. This framework employs a diffusion model with a determinant-based sampling strategy to generate diverse low-level trajectories, improving both planning success rates and task generalization.
@@
In the context of STL trajectory planning, the use of generative models has also been explored recently. For example, Zhong et al. (2023) proposed a classifier-based guidance approach that leverages robustness gradients to guide diffusion model sampling, enabling the generation of vehicle trajectories that adhere to traffic rules specified by STL. Building on this, Meng and Fan (2024) introduced a data augmentation method to enhance trajectory diversity and improve rule satisfaction rates. However, these approaches are still limited to simpler STL tasks, primarily due to the complexity of optimizing robustness values and the inherent trade-off between maximizing reward objectives and maintaining the feasibility of the generated trajectories Li et al. (2024).
@@
In this paper, we address the challenge of generating trajectories for complex, long-horizon STL tasks that are feasible for an underlying system with unknown dynamics. Specifically, we assume access only to a set of task-agnostic trajectory data from previous operations. Inspired by recent advances in decomposition-based STL planningKapoor et al. (2024), we propose a novel hierarchical framework that integrates task decomposition, search algorithms, and generative models. First, complex STL tasks are decomposed into a set of time-aware reach-avoid progresses and time constraints. Next, a search algorithm, heuristically guided by the trajectory data, is employed to allocate these progresses and generate a sequence of waypoints with corresponding timestamps. Finally, a pre-trained diffusion model, trained on task-agnostic data, is used to sequentially generate trajectories that achieve the timed waypoints, resulting in a complete solution. To the best of our knowledge, our algorithm is the first data-driven approach with zero-shot generalization capabilities for complex STL tasks. Simulation experiments demonstrate that our method achieves a high success rate in trajectory planning across diverse, long-horizon STL tasks and outperforms commonly used non-data-driven methods in terms of efficiency.
@@ 2.2 Signal Temporal Logic
We use signal temporal logic (STL) to describe the formal task imposed on the generated state sequence Maler and Nickovic (2004). Specifically, we consider STL formula in the Positive Normal Form (PNF) Sadraddini and Belta (2015) whose syntax is as follows:
@@
where $\top$ is the true predicate and $\mu$ is an atomic predicate associated with an evaluation function $h_{\mu}:\mathbb{R}^{n}\to\mathbb{R}$, i.e., predicate $\mu$ is true at state $\mathbf{x}_{t}$ iff $h_{\mu}(\mathbf{x}_{t})\geq 0$. Furthermore, $\wedge$ and $\vee$ are logic operators “conjunction” and “disjunction”, respectively; $\mathrm{U}_{[a,b]},\mathrm{F}_{[a,b]}$ and $\mathrm{G}_{[a,b]}$ are temporal operators “until”, “eventually” and “always”, respectively; $[a,b]$ is a time interval such that $a,b\in\mathbb{Z},0\leq a\leq b<\infty$. Note that, negation is not used in the PNF. However, as shown in Sadraddini and Belta (2015), this does not result in any loss of generality as one can always redefine atomic predicates to account for the presence of negations, allowing any general STL formula to be expressed in PNF. In our work, we impose an additional restriction on the Prenex Normal Form of formulas. Specifically, for any formula of the form $\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2}$, $\varphi_{1}$ can only involve temporal operator “always”. This restriction is introduced for technical reasons, as it facilitates the decomposition of the overall formula into a set of progresses.
For any signal $\mathbf{s}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$, we denote by $\mathbf{s}_{t}\vDash\varphi$ if $\mathbf{s}$ satisfies STL formula $\varphi$ at time $t$, and we denote by $\mathbf{s}\vDash\varphi$ if $\mathbf{s}_{0}\vDash\varphi$. This is formal defined by the Boolean semantics of STL formulae as follows Bartocci et al. (2018):
$\mathbf{s}_{t}\vDash\mu$ $\Leftrightarrow h_{\mu}(\mathbf{x}_{t})\geq 0,$ (3)
$\mathbf{s}_{t}\vDash\varphi_{1}\wedge\varphi_{2}$ $\Leftrightarrow\mathbf{s}_{t}\vDash\varphi_{1}\wedge\mathbf{s}_{t}\vDash\varphi_{2},$ (4)
$\mathbf{s}_{t}\vDash\varphi_{1}\vee\varphi_{2}$ $\Leftrightarrow\mathbf{s}_{t}\vDash\varphi_{1}\vee\mathbf{s}_{t}\vDash\varphi_{2},$ (5)
$\mathbf{s}_{t}\vDash\mathrm{F}_{[a,b]}\varphi$ $\Leftrightarrow\exists t^{\prime}\in[t+a,t+b]\text{ s.t. }\mathbf{s}_{t^{\prime}}\vDash\varphi,$ (6)
$\mathbf{s}_{t}\vDash\mathrm{G}_{[a,b]}\varphi$ $\Leftrightarrow\forall t^{\prime}\in[t+a,t+b]\text{ s.t. }\mathbf{s}_{t^{\prime}}\vDash\varphi,$ (7)
$\quad\wedge\forall t^{\prime\prime}\in\left[t,t^{\prime}\right],\mathbf{s}_{t^{\prime\prime}}\vDash\varphi_{1}.$ (8)
@@ 2.3 Planning with Unknown Dynamics
In the context of STL planning, the objective is to determine an action sequence such that the resulting signal satisfies the specified STL formula. When the system dynamics are perfectly known, this problem can be solved using model-based optimization approaches (e.g., see Raman et al. (2014); Kurtz and Lin (2022); Sun et al. (2022)). In contrast, our work addresses a setting with unknown dynamics. Specifically, we assume the mapping $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is unknown,but a dataset of historical operational trajectories, consistent with the underlying unknown system dynamics, is available. Note that each trajectory in the dataset is collected from the previous task-agnostic operations and may vary in length. Our goal is to leverage these task-agnostic trajectories to generate new trajectories that satisfy any given STL formula, thereby achieving zero-shot task generalization at test time.
@@
Given a set of trajectories from the unknown system (1) and a STL formula $\varphi$, find a sequence of actions $\mathbf{a}_{0}\mathbf{a}_{1}\dots\mathbf{a}_{T}$ such that the resulting signal $\mathbf{s}$ satisfies the STL formula, i.e., $\mathbf{s}\vDash\varphi$.
@@ 3.1 Overall Framework
Figure 1: The Overall Framework of Our Proposed Method
First, we provide an overview of our proposed planning methods, whose overall structure is illustrated in Figure 1. Specifically, our method consists of the following three parts:
  • Task Decomposition: First, we decompose the given STL into a set of spatial-temporal relevant progresses $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}^{\mathcal{I}}$, where $\mathbb{P}^{\mathcal{R}}$ is the set of reachability progresses and $\mathbb{P}^{\mathcal{I}}$ is the set of invariance progresses. Furthermore, these progresses need to be satisfied subject to a set of time constraints $\mathbb{T}$ over a set of time variables $\Lambda$.
  • Progress Allocation: While the decomposition above is task-focused, achieving progress in the correct order requires consideration of the underlying dynamics of the system. To address this, we employ a pre-trained time predictor derived from trajectory data to estimate the time required for the system to transition from one waypoint to another. A search-based algorithm is then used to determine a sequence of waypoints with associated timestamps $(\mathbf{x}_{0},t_{0})(\mathbf{x}_{1},t_{1})\dots(\mathbf{x}_{n},t_{n})$ such that each waypoint satisfies the corresponding reachability progress in $\mathbb{P}^{\mathcal{R}}$, and all timestamps comply with the time constraints $\mathbb{T}$.
  • Trajectory Generation: The final step is to generate an executable trajectory for the system, ensuring that each waypoint is visited at the correct time while maintaining satisfaction of the invariance progresses. This problem can be framed as several reach-avoid control problems with unknown system dynamics. Our approach leverages a diffusion model, pre-trained on trajectory data, to generate task-compliant and dynamic-feasible trajectories through constraints-guided sampling. Following Ajay et al. (2022), we utilize the diffusion model solely for generating the trajectory’s state sequence. The action sequence can be obtained offline using an inverse dynamics model Agrawal et al. (2016) or online via appropriate controllers.
Nexet, we provide the technical details of each part.
@@ 3.2 Decompositions of STL Formulae
Eliminating Disjunctions
For the given STL task $\varphi$, our first step is to covert it into the disjunctive normal form (DNF) $\tilde{\varphi}=\varphi_{1}\vee\varphi_{2}\vee...\vee\varphi_{n}$, where each subformula $\varphi_{i}$ involves no “disjunction”. Formally, for any STL formula, one can obtain its DNF by recursively applying the following replacements:
  • replace $\mathrm{F}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ by $\mathrm{F}_{[a,b]}\varphi_{1}\vee\mathrm{F}_{[a,b]}\varphi_{2}$;
  • replace $\mathrm{G}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ by $\mathrm{G}_{[a,b]}\varphi_{1}\vee\mathrm{G}_{[a,b]}\varphi_{2}$;
  • replace $(\phi_{1}\!\vee\!\phi_{2})\mathrm{U}_{[a,b]}(\varphi_{1}\!\vee\!\varphi_{2})$ by $\bigvee_{i,j\in\{1,2\}}\phi_{i}\mathrm{U}_{[a,b]}\varphi_{j}$.
@@
Note that the last two replacements are not equivalent, making the resulting DNF $\tilde{\varphi}$ stronger than the original formula ${\varphi}$. Furthermore, to achieve the task defined by $\tilde{\varphi}$, it suffices to satisfy one of the subformulas $\varphi_{i}$. Without loss of generality, we will assume henceforth that the DNF contains only a single subformula, as the STL planning problem can be addressed for each subformula individually. In other words, moving forward, we will focus on STL formulae, denoted directly by $\varphi$, without negations (due to the PNF) and without disjunctions (due to the DNF).
Progresses and Constraints
Next, we further decompose the overall STL task $\varphi$ into a set of progresses. Specifically, we consider the following two types of progresses:
@@
  • Reachability Progress: we denote by $\mathcal{R}(a_{\Lambda},b_{\Lambda},\mu)$ that $\exists t\in[a_{\Lambda},b_{\Lambda}],\mathbf{x}_{t}\vDash\mu$.

### v2

@@ Abstract
Signal Temporal Logic (STL) is a powerful specification language for describing complex temporal behaviors of continuous signals, making it well-suited for high-level robotic task descriptions. However, generating executable plans for STL tasks is challenging, as it requires consideration of the coupling between the task specification and the system dynamics. Existing approaches either follow a model-based setting that explicitly requires knowledge of the system dynamics or adopt a task-oriented data-driven approach to learn plans for specific tasks. In this work, we address the problem of generating executable STL plans for systems with unknown dynamics. We propose a hierarchical planning framework that enables zero-shot generalization to new STL tasks by leveraging only task-agnostic trajectory data during offline training. The framework consists of three key components: (i) decomposing the STL specification into several progresses and time constraints, (ii) searching for timed waypoints that satisfy all progresses under time constraints, and (iii) generating trajectory segments using a pre-trained diffusion model and stitching them into complete trajectories. We formally prove that our method guarantees STL satisfaction, and simulation results demonstrate its effectiveness in generating dynamically feasible trajectories across diverse long-horizon STL tasks. Project Page: https://cps-sjtu.github.io/Zero-Shot-STL/
@@ 1 Introduction
Signal Temporal Logic (STL) is a formal specification language used to describe the temporal behavior of continuous signals. It has become widely adopted for specifying high-level robotic behaviors due to its expressiveness and the availability of both Boolean and quantitative evaluation measures. Controlling robots under STL task constraints, however, is a challenging problem, as it requires balancing both the satisfaction of the task and the feasibility of the system dynamics. In cases where the environment and system dynamics are fully known, several representative methods have been developed, including optimization-based approaches [1, 2, 3], gradient-based techniques [4, 5], and sampling-based methods [6]. However, these methods are often difficult to apply in practical scenarios, where the system dynamics and environment are either unknown or difficult to model.
@@
To address the challenge of unknown dynamics, several learning-based approaches have been proposed. One typical method is reinforcement learning (RL) [7, 8, 9, 10, 11, 12], where an appropriate reward function is designed to approximate the satisfaction of the STL task. However, these methods often struggle with long-horizon STL tasks and lack generalization capabilities across different tasks. Another approach involves first learning a system model and then integrating it with model-based planning methods. For example, in [13], the authors trained a neural network to approximate the system dynamics and combined it with an optimization-based approach. However, this method is limited to simple short-horizon STL tasks due to its high computational cost. In [14], the authors used goal-conditioned RL to train multiple goal-conditioned policies, referred to as “skills," to accomplish specific objectives. They then applied a search algorithm to determine the optimal sequence of “skills" needed to satisfy the given STL tasks. While this approach enables a certain degree of task generalizations, these tasks must be based on pre-defined objectives associated with the skills.
@@
More recently, generative models, such as diffusion models [15], have emerged as a new approach for generating trajectories for systems with unknown dynamics [16, 17, 18, 19, 20], gaining popularity across various applications. Compared to traditional model-based reinforcement learning methods, these generative approaches are better suited for long-horizon decision-making and offer greater test-time flexibility [16], making them particularly effective for complex tasks. For example, for finite Linear Temporal Logic ($\text{LTL}_{f}$) tasks, [21] introduced a classifier-based guidance approach to steer the sampling of diffusion models, ensuring that generated trajectories satisfy $\text{LTL}_{f}$ requirements. Similarly, [22] proposed a hierarchical framework that decomposes co-safe LTL tasks into sub-tasks using hierarchical reinforcement learning. This framework employs a diffusion model with a determinant-based sampling strategy to generate diverse low-level trajectories, improving both planning success rates and task generalization.
@@
In the context of STL trajectory planning, the use of generative models has also been explored recently. For example, [23] proposed a classifier-based guidance approach that leverages robustness gradients to guide diffusion model sampling, enabling the generation of vehicle trajectories that adhere to traffic rules specified by STL. Building on this, [24] introduced a data augmentation method to enhance trajectory diversity and improve rule satisfaction rates. However, these approaches are still limited to simpler STL tasks, primarily due to the complexity of optimizing robustness values and the inherent trade-off between maximizing reward objectives and maintaining the feasibility of the generated trajectories [25].
@@
In this paper, considering that trajectories satisfying complex STL specifications are typically difficult to collect in real-world scenarios, we focus on composing such trajectories by stitching together short trajectory chunks. The main challenge lies in determining appropriate ways to combine these chunks such that the resulting trajectory satisfies the global STL specification while maintaining dynamic consistency and feasibility. Inspired by recent advances in decomposition-based STL planning[26], we propose a novel hierarchical framework that integrates task decomposition, search algorithms, and generative models. First, complex STL tasks are decomposed into a set of time-aware reach-avoid progresses. Next, a search algorithm, heuristically guided by the trajectory data, is employed to allocate these progresses and generate a sequence of waypoints with corresponding timestamps. Finally, a pre-trained diffusion model, trained on task-agnostic data, is used to sequentially generate trajectory chunks that connect adjacent timed waypoints. All trajectory chunks are then stitched together to form the complete trajectory.
To the best of our knowledge, our algorithm is the first data-driven approach with zero-shot generalization capability for complex STL tasks. We have formally proven the soundness of our STL decomposition and planning algorithm, which guarantees that the generated trajectories satisfy any given STL specification. Furthermore, we empirically evaluate the dynamic consistency and feasibility of the planned trajectories through simulation experiments. Simulation results demonstrate that our method achieves a high execution success rate across diverse long-horizon STL tasks, where the diffusion-based baseline fails. Moreover, it outperforms both the diffusion-based baseline and a standard non-data-driven method in planning efficiency.
@@ 2.2 Signal Temporal Logic
We use signal temporal logic (STL) to describe the formal task imposed on the generated state sequence [27]. Specifically, we consider STL formula in the Positive Normal Form (PNF) [28] whose syntax is as follows:
@@
where $\top$ is the true predicate and $\mu$ is an atomic predicate associated with an evaluation function $h_{\mu}:\mathbb{R}^{n}\to\mathbb{R}$, i.e., predicate $\mu$ is true at state $\mathbf{x}_{t}$ iff $h_{\mu}(\mathbf{x}_{t})\geq 0$. Furthermore, $\wedge$ and $\vee$ are logic operators “conjunction" and “disjunction", respectively; $\mathrm{U}_{[a,b]},\mathrm{F}_{[a,b]}$ and $\mathrm{G}_{[a,b]}$ are temporal operators “until", “eventually" and “always", respectively; $[a,b]$ is a time interval such that $a,b\in\mathbb{Z},0\leq a\leq b<\infty$. Note that, negation is not used in the PNF. However, as shown in [28], this does not result in any loss of generality as one can always redefine atomic predicates to account for the presence of negations, allowing any general STL formula to be expressed in PNF. In our work, we impose an additional restriction on the Prenex Normal Form of formulas. Specifically, for any formula of the form $\varphi_{1}\mathrm{U}_{[a,b]}\varphi_{2}$, $\varphi_{1}$ can only involve temporal operator “always”. This restriction is introduced for technical reasons, as it facilitates the decomposition of the overall formula into a set of progresses. For any signal $\mathbf{s}=\mathbf{x}_{0}\mathbf{x}_{1}\dots\mathbf{x}_{T}$, we denote by $\mathbf{s}_{t}\vDash\varphi$ if $\mathbf{s}$ satisfies STL formula $\varphi$ at time $t$, and we denote by $\mathbf{s}\vDash\varphi$ if $\mathbf{s}_{0}\vDash\varphi$. This is formally defined by the Boolean semantics of STL formulae as shown in Appendix C.1.
@@ 2.3 Planning with Unknown Dynamics
In the context of STL planning, the objective is to determine an action sequence such that the resulting signal satisfies the specified STL formula. When the system dynamics are perfectly known, this problem can be solved using model-based optimization approaches (e.g., see [1, 2, 3]). In contrast, our work addresses a setting with unknown dynamics. Specifically, we assume the mapping $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ is unknown,but a dataset of historical operational trajectories, consistent with the underlying unknown system dynamics, is available. Note that each trajectory in the dataset is collected from the previous task-agnostic operations and may vary in length. Our goal is to leverage these task-agnostic trajectories to generate new trajectories that satisfy any given STL formula, thereby achieving zero-shot task generalization at test time.
@@
Given a set of trajectories from the unknown system (1) and a STL formula $\varphi$ in the desired form, find a sequence of actions $\mathbf{a}_{0}\mathbf{a}_{1}\dots\mathbf{a}_{T}$ such that the resulting signal $\mathbf{s}$ satisfies the STL formula, i.e., $\mathbf{s}\vDash\varphi$.
@@ 3.1 Overall Framework
First, we provide an overview of our proposed planning framework, whose overall structure is illustrated in Figure 1. Our method consists of three key components. The first is semantics-based task decomposition, in which the given STL formula is decomposed into a set of spatial-temporal progresses $\mathbb{P}=\mathbb{P}^{\mathcal{R}}\dot{\cup}\mathbb{P}$$^{\mathcal{I}}$, where $\mathbb{P}^{\mathcal{R}}$ denotes the set of reachability progresses and $\mathbb{P}^{\mathcal{I}}$ the set of invariance progresses. These progresses must be satisfied under a set of time constraints $\mathbb{T}$ defined over time variables $\Lambda$. The second component is dynamics-aware progress allocation. While the above decomposition is task-centric, ensuring that the progresses can be satisfied in the correct temporal order requires considering the system’s underlying dynamics. To this end, we use a pre-trained Time Predictor, learned from task-agnostic trajectory data, to estimate the time required to transition between waypoints. A search-based algorithm then determines a sequence of timed waypoints $(\tilde{\mathbf{x}}_{0},t_{0}),(\tilde{\mathbf{x}}_{1},t_{1}),\dots,(\tilde{\mathbf{x}}_{n},t_{n})$ such that each waypoint satisfies a corresponding reachability progress in $\mathbb{P}^{\mathcal{R}}$ and complies with all invariance progresses in $\mathbb{P}^{\mathcal{I}}$.
Figure 1: The Overall Framework of Our Proposed Method.
The third component is trajectory generation. Given the sequence of timed waypoints, we generate an executable trajectory that ensures each waypoint is visited at the correct time while satisfying all invariance progresses by stitching together shorter trajectory segments. We instantiate this using existing constrained sampling frameworks (e.g., SafeDiffuser [29]), which leverage diffusion models pre-trained on trajectory data to produce dynamically feasible, constraint-compliant trajectories through constrained sampling. Next, we provide the technical details of each component.
@@ 3.2 Decompositions of STL Formulae
Eliminating Disjunctions. For the given STL task $\varphi$, our first step is to covert it into the disjunctive normal form (DNF) $\tilde{\varphi}=\varphi_{1}\vee\varphi_{2}\vee\cdots\vee\varphi_{n}$, where each subformula $\varphi_{i}$ involves no “disjunction". Formally, for any STL formula, one can obtain its disjunctive normal form (DNF) by recursively applying the following replacements: (i) replace $\mathrm{F}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ with $\mathrm{F}_{[a,b]}\varphi_{1}\vee\mathrm{F}_{[a,b]}\varphi_{2}$; (ii) replace $\mathrm{G}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ with $\mathrm{G}_{[a,b]}\varphi_{1}\vee\mathrm{G}_{[a,b]}\varphi_{2}$; and (iii) replace $(\phi_{1}\vee\phi_{2})\mathrm{U}_{[a,b]}(\varphi_{1}\vee\varphi_{2})$ with $\bigvee_{i,j\in\{1,2\}}\phi_{i}\mathrm{U}_{[a,b]}\varphi_{j}$. Note that the last two replacements are not equivalent, making the resulting DNF $\tilde{\varphi}$ stronger than the original formula $\varphi$}; i.e., $\tilde{\varphi}\Rightarrow\varphi$ but not conversely, adding conservativeness while preserving soundness. Furthermore, to achieve the task defined by $\tilde{\varphi}$, it suffices to satisfy one of the subformulas $\varphi_{i}$. Without loss of generality, we will assume henceforth that the DNF contains only a single subformula, as the STL planning problem can be addressed for each subformula individually. In other words, moving forward, we will focus on STL formulae, denoted directly by $\varphi$, without negations (due to the PNF) and without disjunctions (due to the DNF). Similar simplifications and assumptions are common in STL planning or control synthesis literature [30].
Progresses. We encode the STL task $\varphi$ by a finite set of progresses together with a set of time-variable constraints. For a signal fragment $\mathbf{s}_{t}=\mathbf{x}_{t}\mathbf{x}_{t+1}\dots\mathbf{x}_{T}$, we use two canonical progress types:
@@
  • Reachability progress $\mathcal{R}(a_{\Lambda}+t,b_{\Lambda}+t,\mu)$: there exists $t^{\prime}\in[a_{\Lambda}+t,b_{\Lambda}+t]$ such that $\mathbf{x}_{t^{\prime}}\vDash\mu$.